1. (Sections 2.3, 2.10, 2.11, 2.12) Let L1 be the line in R 3 with equation (x, y, z) = (1, 0, 2) + t(−1, 3, 4) ; t ∈ R and let L2 be the line that is parallel to L1 and contains the point (1, −1, 3). Let V be the plane that contains both the lines L1 and L2. (a) Find two vectors that are both parallel to the plane V but are not parallel to one another. (2) (b) Find a vector that is perpendicular to the plane V . (2) (c) Find an equation for the plane V . (2) (d) Find an equation for the line L3 that is perpendicular to the plane V and contains the point (1, −1, 4) . (2) Hint: Find a parametric equation for L3. Don’t try to find a Cartesian equation for L3. (Study Remarks 2.12.2.)


1
Expert's answer
2021-04-27T01:47:09-0400

"L_1 : (x,y,z)=(1,0,2)+t(-1,3,4), t\in R\\
L_1 ||L_2 \\
L_1 : (x,y,z)=(1,-1,3)+t(-1,3,4), t\in R\\
V: L_1\in V, L_2\in V"

a)

"\vec{a}=(-1,3,4)||V\\
A(1,0,2)\in L_1\\
B(1,-1,3)\in L_2\\
\vec{AB}=(1-1,-1-0,3-2)=(0,-1,1)|| V\\
\vec{a}\not|| \vec{AB}\\
\frac{-1}{0}\not=\frac{3}{-1}\not=\frac{4}{1}"

b)

"\vec{n}=[\vec{a},\vec{AB}]=\begin{vmatrix}
 \vec{i} & \vec{j}&\vec{k} \\
 -1&3&4\\
0&-1&1 
\end{vmatrix}=\\
=\vec{i}(3+4)-\vec{j}(-1-0)+\vec{k}(1-0)=(7,1,1)"

"\vec{n}" is perpendicular to the plane "V"

c)

"\vec{n}=(7,1,1)" is perpendicular to the plane, "A(1,0,2)\in V"

"7(x-1)+1(y-0)+1(z-2)=0\\
7x+y+z-9=0"

d)

"L_3" is perpendicular to the plane "V"

"\vec{n}=(7,1,1)||L_3\\
C(1,-1,4)\in L_3\\
(x,y,z)=(1,-1,4)+t(7,1,1), t\in R"